OpenStudy (anonymous):

Find the first 5 terms of the sequence (explicit): a n = 3n-1

jimthompson5910 (jim_thompson5910):

First term: an = 3n - 1 a1 = 3(1) - 1 a1 = 3 - 1 a1 = 2 So the first term is a1 = 2 ------------------------------------------------------- Second term: an = 3n - 1 a2 = 3(2) - 1 a2 = 6 - 1 a2 = 5 So the first term is a2 = 5 ------------------------------------------------------- I'll let you finish and find the next 3 terms

jimthompson5910 (jim_thompson5910):

oops meant to say "So the second term is a2 = 5"